Suppose the inflation rate will be 4.5% for the next 12 years. What will a dozen large brown eggs cost in 9 years if it costs $2.25 toda
Answer by addingup(3677)

2.25(1+0.045)^9=
2.25(1.045)^9=
2.25(1.4861)= 3.34